# Magnetically charged black-bounce solution via nonlinear electrodynamics in a k-essence theory

### v1

@@ Abstract
In the present work, we obtain and analyze a new class of analytical solutions of magnetically charged black bounces in k-essence theory, spherically symmetric in (3+1)-dimensions, coupled to nonlinear electrodynamics (NED). We consider two metric models, Simpson-Visser and Bardeen, for the k-essence configurations $n=1/3$ and $n=1/5$. We obtain in an analytical way which scalar field, field potential, and Lagrangian NED are necessary to support the metrics. We analyze the behavior of these quantities and the energy conditions due to the scalar field and the NED.
@@ I Introduction
General relativity in its classical form is the simplest theory to describe gravitational and cosmological scenarios. Several objects can be predicted within this framework, including black holes, wormholes, neutron stars, and interesting phenomena such as gravitational waves and light deviation in the presence of a strong field, among others [1, 2, 3, 4, 5, 6]. The best-known spherically symmetric black hole solution is the Schwarzschild solution. This solution is quite simple, as it is described solely by the mass of the black hole and has only one event horizon, which is a non-return surface, and one physical singularity, a point where geodesics are interrupted. Regular black holes emerge as alternatives that stand out compared to typical black holes due to the absence of singularities within their interiors, with the first being proposed by Bardeen in 1968 [7]. Later, Beato and Garcia showed that the matter content associated with the Bardeen metric that solved Einstein’s equations was nonlinear electrodynamics [8]. Usually, static black holes with regularized centers can always have nonlinear electrodynamics as their source of matter, with the sources being either electric or magnetic [9, 10, 12, 11]. The same is not necessarily true for rotating solutions [13].
@@
Recently, Simpson and Visser used the Schwarzschild metric with a regularization procedure, $r^{2}\to r^{2}+a^{2}$, where $a$ is the regularization parameter, to obtain the so-called black-bounce solutions [14]. This regularization method essentially removes the point where the singularity existed, along with its surrounding neighborhood, thereby creating a sort of wormhole within a black hole. These spacetimes can be classified as two-way traversable wormholes ($a>2m$), one-way wormholes ($a=2m$), and regular black holes with symmetric horizons ($a<2m$), where $a$ represents the throat of the wormhole. In the case $a=2m$, the event horizon coincides with the position of the wormhole’s throat, and in this situation, we have what is called a black throat. In works [15, 16], it was verified that the matter content associated with the Simpson-Visser metric is a combination of nonlinear electrodynamics and a phantom scalar field, where the electromagnetic source is a magnetic charge. It is important to note that these same solutions can also be obtained by considering an electrically charged source [17].
@@
Following Simpson and Visser’s work, a wide range of scenarios have been explored. These include modifications in the area functions [18], modifications in the mass of the object making it depend on the radial coordinate [19], and a regularization of the Reissner–Nordström spacetime [20, 21]. Furthermore, there are works in the context of modified gravity [22, 23, 24], non-conservative theory [25], Braneworld [26], and black strings [27, 28, 29]. This class of solutions has also been extended to spherically symmetric and stationary spacetimes [20, 30, 28, 31], as well as to the study of light deflection and gravitational lensing effects [32, 33, 34, 35, 36]. In the mass limit tending to zero, the Simpson-Visser spacetime transforms into the topologically charged Ellis-Bronnikov spacetime, and in this context, quantum systems were explored [37, 38, 39, 40, 41, 42].
@@
Conventionally, in general relativity, solutions for regular black holes and traversable wormholes require that energy conditions be violated, thus requiring exotic matter. For the case of regular black holes with only nonlinear electrodynamics, the strong energy condition is always violated. However, there are also solutions that violate other energy conditions [43]. In the case of the scalar field, a canonical scalar field minimally coupled with gravitational theory is not sufficient to generate regular solutions, necessitating the presence of a phantom scalar field, which violates the null energy condition [44].
@@
The initial proposal of this work was to investigate possible modifications in the energy conditions, for example, arising from nonlinear electrodynamics incorporated into a k-essence theory. As we will see below, the electromagnetic sector is not modified by the presence of the k-essence function, however, the phantom and potential scalar field configurations are altered.
@@
The paper is structured as follows: Section II establishes the theoretical basis of the k-essence model in conjunction with nonlinear electrodynamics, including the derivation of the equations of motion. In Sections III and IV, we apply the methodology to two specific models and derive the physical quantities for the k-essence configurations of $n=1/3$ and $n=1/5$. In Section V, we derive the energy conditions generically for each configuration, $n=1/3$ and $n=1/5$. In Section VI, we analyze the power conditions for each model. Finally, the conclusions are presented in Section VII.
@@ II General relations
The k-essence theories are characterized by the presence of scalar fields whose kinetic terms are introduced in a non-canonical way and have been recently explored in the context of wormhole solutions [45, 46]. Now, we will introduce a term representing nonlinear electrodynamics. Therefore, consider the model described by the action below:
@@
By varying the above action 1 with respect to the fields and the metric tensor, we obtain the following equations of motion:
@@
where the electromagnetic field is defined by
@@
Thus, we write the general equations of motion, which are the same as those contained in Refs.[45, 46, 51]. However, they are now modified by nonlinear electrodynamics. It is assumed that the function $X=-\eta e^{-2\alpha}(\phi^{\prime})^{2}$ is positive, which implies that $\eta=-1$. As a result, the equations of motion take the form:
@@
It has been established in previous works [45, 46] that pursuing black-bounce solutions solely with the kinetic term of the k-essence function is not mathematically consistent. Therefore, when constructing these new solutions, we must incorporate a scalar potential given by $F(X)=F_{0}X^{n}-2V(\phi)$, where $F_{0}$ is a constant, $n$ is a real number, and $V(\phi)$ is the potential function. Using Eq. (15), we can derive a general expression for the scalar field that depends on both the angular metric function $\Sigma(x)$ and the metric function $A(x)$, unlike most studies, where scalar field depends solely on the angular function and its derivatives [51, 44, 52].
@@ III First Model
In this section, we will consider the ingredients of the black-bounce solution proposed by Simpson-Visser in Ref. [14]. Specifically, we assume that the throat radius is exactly the magnetic charge, $a=q_{m}$. Therefore, the metric functions are given by:
@@
In the subsection below, we will calculate all quantities present in the equations of motion (15-18) using metric functions Eq. (19).
@@
$\phi(x)=\left(\frac{6q^{2}_{m}}{F_{0}}\right)^{3/2}\left[\frac{x}{4q^{2}_{m}{\Sigma}^{4}}+\frac{3x}{8q^{4}_{m}{\Sigma}^{2}}-\frac{2mx(15q^{4}_{m}+8x^{4}+20x^{2}{q^{2}_{m}})}{15q^{6}_{m}{\Sigma}^{5}}+\frac{3\arctan\left(\frac{x}{q_{m}}\right)}{8q^{5}_{m}}\right],$ (20)
@@
Finally, using Eqs. (17)-(18), we can find the scalar potential $V(x)$ as well as the electromagnetic quantity $L(f)$. Therefore, we have:
@@
$\phi\left(x\to{\infty}\right)=-\phi\left(x\to{-\infty}\right)=-\frac{\sqrt{3/2}\left(45\pi{q_{m}}-256m\right)}{20q^{3}_{m}},$ (27)
@@
Figure 1: The scalar field, given by Eq. (20), is represented in a) for different charge values and considering regions inside and outside any possible horizon. Similarly, we have the potential, given by Eq. (21), in b), and the electromagnetic functions, given by Eq. (22) in c) and Eq. (23) in d). All these are for $n=1/3$ setting, and we are adopting $F_{0}=1$.
In panel 1, we have the graphical representation of the physical quantities obtained in the model above for the configuration of $n=1/3$. Regarding the scalar field (1), looking at its asymptotic form Eq. (27), it is clear that there is a limit for charge values at which sign inversion occurs. For instance, for $x\to{+\infty}$, the field is positive for charge values around $q_{m}\approx{1.81m}$. Regarding the potential (1), we can verify the same behavior presented in works [45, 46], where ,for certain charge values, already within the event horizon, the curves begin to create a minimum that grows as the charge becomes more internal. Likewise, we have the behavior of the electromagnetic functions in figures (1) and (1).
@@
An observation to be made is that the electromagnetic quantity $L_{f}$, obtained by Eq. (15), does not depend on the scalar field and is therefore the same as that obtained in the previous section III.1. Likewise, we can write the electromagnetic quantity $L(f)$ in terms of the invariant as done in the previous section. Therefore, we have
@@
We can also express the scalar potential for this field configuration using the most appropriate variables as done in the previous section. Thus, we have
@@
Figure 2: Scalar field, Eq. (28), in panel a) and potential, Eq. (29), in panel b) for different charge values and considering regions inside and outside any possible horizon, for the configuration $n=1/5$. Here, we define $F_{0}=1$.
In panel 2, we have the graphical representation of the scalar field, Eq. (28), and potential, Eq. (29), for the configuration $n=1/5$. As illustrated in figure 2, in the limit of $x\to\infty$ the scalar field grows for increasingly larger charge values and tends to zero for small values of the radial coordinate. The behavior of the scalar potential is qualitatively the same as the previous configuration (III.1), changing only by a numerical shift (Figure 2).
@@ IV Second Model
As discussed in Refs. [52, 19], this space-time has four horizons, two event horizons and two Cauchy horizons. The extreme horizon is obtained for $q_{m}=q_{ext}=4m/3\sqrt{3}$ and the others are symmetric and labeled by $(-x_{+},-x_{C},x_{C},x_{+})$. Where $x_{+}$ is the event horizon and $x_{C}$ is the Cauchy horizon.
@@
Following the same steps carried out in section III.1, we can obtain all physical quantities of interest for this space-time. Therefore, we have
@@
Figure 3: Scalar field Eq. (34) for some charge values and considering regions inside and outside any possible horizon is represented in a). Similarly, we have the potential Eq. (36) b) and the electromagnetic functions Eq. (35) c) and Eq. (37) d). All for $n=1/3$ setting, adopting $F_{0}=1$.
In panel 3, we have a graphical representation of the physical quantities obtained above for the configuration $n=1/3$ for Bardeen-type spacetime. In figure 3, we show the behavior of the scalar field, Eq. (34), in all regions of the spacetime, considering different charge values. For some values of charge, $q_{m}>q_{ext}=\frac{4m}{3\sqrt{3}}$, there is no event horizon. Note in the asymptotic form of the scalar field Eq. (40) that for a certain limiting value of charge within the horizon, around $q_{m}\approx{0.26m}$, the scalar field tends to reverse sign.
For the scalar potential 3, we observe a behavior similar to the cases previously investigated. For regions outside the external horizon, the potential behaves like a barrier. For regions within the horizon, the potential starts to exhibit two symmetrical minima that tend to grow as the charge becomes more internal to extreme radius. The electromagnetic quantities Eq. (35) and Eq. (37) exhibit behavior similar to that presented in section III.1 and can be visualized in figures 3 and 3.
@@
Figure 4: Scalar field, Eq. (41), in panel a) and potential, Eq. (42), in panel b) with different charge values and considering regions inside and outside any possible horizon, for the configuration $n=1/5$ Here, we define $F_{0}=1$.
In panel 4, we have the graphical representation for the scalar field Eq. (41) and potential Eq. (42) with a configuration $n=1/5$. Qualitatively, the scalar field 4 and the potential 4 exhibit similar behavior to the previously investigated cases, with only minor numerical shifts.
@@
The same analysis can be extended to the second model (Bardeen type) when comparing the functions Eqs. (38) and (45) as well as the expression Eq. (35) which are the same found for the canonical scalar field [52].
@@ V General Energy Conditions
We know that black bounce solutions violate the null energy condition, which must violate all other energy conditions [14, 53]. In this section, we will make the generic construction for the energy conditions, that is, just considering the static and spherically symmetric metric Eq. (14), as well as the stress-energy tensor defined in Eqs. (5-6). The energy conditions are defined by the combination of energy density and radial and tangential pressures as [^1] In the notation we follow, NEC is used for the null energy condition, WEC for the weak energy condition, SEC for the strong energy condition, and DEC for the dominant energy condition. The subscript $1$ is used when we have the energy density with radial pressure, $2$ when we have the energy density with tangential pressure and $3$ when we have both pressures or just the energy density.
@@
where the energy density is denoted by $\rho^{\phi,EM}$, while $p^{\phi,EM}_{1}$ and $p^{\phi,EM}_{2}$ represent the radial and tangential pressures, respectively. This expression for the stress-energy tensor above Eq. (50) refers to the part outside the event horizon $A>0$. Within the event $A<0$ horizon we have:

### v2

@@ Abstract
In the present work, we obtain and analyze a new class of analytical solutions of magnetically charged black-bounces in the k-essence theory, spherically symmetric in (3+1)-dimensions, coupled to nonlinear electrodynamics (NED). We consider two metric models, Simpson-Visser and Bardeen-type black-bounces, for the k-essence configurations $n=1/3$ and $n=1/5$. We obtain in an analytical way the scalar field, the field potential, and Lagrangian NED, which are necessary to support the metrics. We analyze the behavior of these quantities and the energy conditions due to the scalar field and the NED.
@@ I Introduction
General relativity in its classical form is the simplest theory to describe gravitational and cosmological scenarios. Several objects can be predicted within this framework, including black holes, wormholes, neutron stars, and interesting phenomena such as gravitational waves and light deviation in the presence of a strong field, among others [1, 2, 3, 4, 5, 6]. The best-known spherically symmetric black hole solution is the Schwarzschild solution. This solution is quite simple, as it is described solely by the mass of the black hole and has only one event horizon, which is a non-return surface, and one physical singularity, a point where geodesics are interrupted. Regular black holes emerge as alternatives that stand out compared to typical black holes due to the absence of singularities within their interiors, the first being proposed by Bardeen in 1968 [7]. Later, Beato and Garcia showed that the matter content associated with the Bardeen metric that solved Einstein’s equations was NED [8]. Usually, static black holes with regularized centers can always have NED as their source of matter, the sources being either electric or magnetic [9, 10, 12, 11]. The same is not necessarily true for rotating solutions [13].
@@
Recently, Simpson and Visser used the Schwarzschild metric with a regularization procedure, $r^{2}\to r^{2}+a^{2}$, where $a$ is the regularization parameter, to obtain the so-called black-bounce solutions [14]. This regularization method essentially removes the point where the singularity existed along with its surrounding neighborhood, thereby creating a sort of wormhole within a black hole. These spacetimes can be classified as two-way traversable wormholes ($a>2m$), one-way wormholes ($a=2m$), and regular black holes with symmetric horizons ($a<2m$), where $a$ represents the throat of the wormhole. In the case $a=2m$, the event horizon coincides with the position of the wormhole’s throat, and in this situation, we have what is called a black throat. In works [15, 16], it was verified that the matter content associated with the Simpson-Visser metric is a combination of NED and a phantom scalar field, where the electromagnetic source is a magnetic charge. It is important to note that these same solutions can also be obtained by considering an electrically charged source [17].
@@
Following Simpson and Visser’s work, a wide range of scenarios have been explored. These include modifications to area functions [18], modifications to the mass of the object that makes it depend on the radial coordinate [19], and a regularization of the Reissner–Nordström spacetime [20, 21]. Furthermore, there are works in the context of modified gravity [22, 23, 24], non-conservative theory [25], Braneworld [26], and black strings [27, 28, 29]. This class of solutions has also been extended to spherically symmetric and stationary spacetimes [20, 30, 28, 31], as well as to the study of light deflection and gravitational lensing effects [32, 33, 34, 35, 36]. In the mass limit tending to zero, the Simpson-Visser spacetime transforms into the topologically charged Ellis-Bronnikov spacetime, and in this context quantum systems were explored [37, 38, 39, 40, 41, 42].
@@
Conventionally, in general relativity, solutions for regular black holes and traversable wormholes require that energy conditions be violated, thus requiring exotic matter. For the case of regular black holes with only NED, the strong energy condition is always violated. However, there are also solutions that violate other energy conditions [43]. In the case of the scalar field, a canonical scalar field minimally coupled with gravitational theory is not sufficient to generate regular solutions, necessitating the presence of a phantom scalar field, which violates the null energy condition [44].
@@
The initial proposal of this work was to investigate possible modifications in the energy conditions, for example, arising from NED incorporated into a k-essence theory. As we will see below, the electromagnetic sector is not modified by the presence of the k-essence function, however, the phantom and potential scalar field configurations are altered.
@@
The paper is structured as follows: Section II establishes the theoretical basis of the k-essence model in conjunction with NED, including the derivation of the equations of motion. In Sections III and IV, we apply the methodology to two specific models and derive the physical quantities for the k-essence configurations of $n=1/3$ and $n=1/5$. In Section V, we derive the energy conditions generically for each configuration, $n=1/3$ and $n=1/5$. In Section VI, we analyze the energy conditions for each model. Finally, the conclusions are presented in Section VII.
@@ II General relations
The k-essence theories are characterized by the presence of scalar fields whose kinetic terms are introduced in a non-canonical way and have been recently explored in the context of wormhole solutions [45, 46]. Now, we will introduce a term that represents NED. Therefore, consider the model described by the action below:
@@
By varying the above action (1) with respect to the fields and the metric tensor, we obtain the following equations of motion:
@@
where the electromagnetic scalar is defined by
@@
Thus, we write the general equations of motion, which are the same as those contained in Refs. [45, 46, 51]. However, they are now modified by NED. It is assumed that the function $X=-\eta e^{-2\alpha}(\phi^{\prime})^{2}$ is positive, which implies that $\eta=-1$. As a result, the equations of motion take the form:
@@
It has been established in previous works [45, 46] that pursuing black-bounce solutions solely with the kinetic term of the k-essence function is not mathematically consistent. Therefore, when constructing these new solutions, we must incorporate a scalar potential given by $F(X)=F_{0}X^{n}-2V(\phi)$, where $F_{0}$ is a constant, $n$ is a real number and $V(\phi)$ is the potential function. Using Eq. (15), we can derive a general expression for the scalar field that depends on both the angular metric function $\Sigma(x)$ and the metric function $A(x)$, unlike most studies, where the scalar field depends solely on the angular function and its derivatives [51, 44, 52].
@@ III First Model
In this section, we will consider the ingredients of the black-bounce solution proposed by Simpson-Visser in Ref. [14]. Specifically, we assume that the radius of the throat is exactly the magnetic charge $a=q_{m}$. Therefore, the metric functions are given by:
@@
In the following subsection, we will calculate all quantities present in the equations of motion (15-18) using metric functions Eq. (19).
@@
$\phi(x)=\left(\frac{6q^{2}_{m}}{F_{0}}\right)^{3/2}\left[\frac{x}{4q^{2}_{m}{\Sigma}^{4}}+\frac{3x}{8q^{4}_{m}{\Sigma}^{2}}-\frac{2mx(15q^{4}_{m}+8x^{4}+20x^{2}{q^{2}_{m}})}{15q^{6}_{m}{\Sigma}^{5}}+\frac{3\arctan\left(\frac{x}{q_{m}}\right)}{8q^{5}_{m}}\right].$ (20)
@@
Finally, using Eqs. (17)-(18), we can find the scalar potential $V(x)$ as well as the electromagnetic quantity $L(f)$. Therefore, we have the following.
@@
$\phi\left(x\to{\infty}\right)=-\phi\left(x\to{-\infty}\right)=-\frac{\sqrt{3/2}\left(256m-45\pi{q_{m}}\right)}{20q^{3}_{m}},$ (27)
@@
Figure 1: The scalar field, given by Eq. (20), is represented in a) for different charge values and considering regions inside and outside any possible horizon. Similarly, we have the potential, given by Eq. (21), in b), and the electromagnetic functions, given by Eq. (22) in c) and Eq. (23) in d). All these are for $n=1/3$ setting, and we are adopting $F_{0}=1$. The vertical lines represent the positions of the event horizons. In the cases presented, horizons exist only for $q_{m}=1.5m$ and $q_{m}=1.8m$, because for $q_{m}>2m$ there are no event horizons.
In Fig. 1, we have the graphical representation of the physical quantities obtained in the model above for the configuration of $n=1/3$. Regarding the scalar field (Fig. 1), looking at its asymptotic form Eq. (27), it is clear that there is a limit for charge values at which sign inversion occurs. For instance, for $x\to{+\infty}$, the field is positive for charge values around $q_{m}\approx{1.81m}$. Regarding the potential (Fig. 1), we can verify the same behavior presented in works [45, 46], where, for certain charge values, already within the event horizon, the curves begin to create a minimum that grows as the charge becomes more internal. Likewise, we have the behavior of the electromagnetic functions in figures (Fig. 1) and (Fig. 1).
@@
An observation to be made is that the electromagnetic quantity $L_{f}$, obtained by Eq. (15), does not depend on the scalar field and is therefore the same as that obtained in the previous Section III.1. Likewise, we can write the electromagnetic quantity $L(f)$ in terms of the invariant, as done in the previous section. Therefore, we have
@@
We can also express the scalar potential for this field configuration using the most appropriate variables, as done in the previous section. Thus, we have
@@
The asymptotic form of the scalar field Eq. (28) is written as:
$\phi\left(x\to{\infty}\right)=-\phi\left(x\to{-\infty}\right)=\frac{5\sqrt{5/2}\left(39690\pi m^{2}-131072mq_{m}+11025\pi q_{m}^{2}\right)}{2016q_{m}^{6}},$ (33)
where $F_{0}=1$.
Figure 2: Scalar field, Eq. (28), in panel a) and potential, Eq. (29), in panel b) for different charge values and considering regions inside and outside any possible horizon, for the configuration $n=1/5$. Here, we define $F_{0}=1$. The vertical lines represent the positions of the event horizons. In the cases presented, horizons exist only for $q_{m}=1.5m$ and $q_{m}=1.6m$, because for $q_{m}>2m$ there are no event horizons.
In Fig. 2, we have the graphical representation of the scalar field, Eq. (28), and potential, Eq. (29), for the configuration $n=1/5$. As illustrated in Fig. 2, in the limit of $x\to\infty$ the scalar field grows for increasingly larger charge values and tends to zero for small values of the radial coordinate. The behavior of the scalar potential is qualitatively the same as the previous configuration (Section III.1), changing only by a numerical shift (Fig. 2). From the asymptotic form for the scalar field in Eq. (33), we find that there are no real values for the charge that allow a sign change in this expression, contrary to what happens in the case $n=1/3$.
@@ IV Second Model
As discussed in Refs. [52, 19], this spacetime has four horizons, two event horizons and two Cauchy horizons. The extreme horizon is obtained for $q_{m}=q_{ext}=4m/3\sqrt{3}$ and the others are symmetric and labeled by $(-x_{+},-x_{C},x_{C},x_{+})$, where $x_{+}$ is the event horizon and $x_{C}$ is the Cauchy horizon.
@@
Following the same steps carried out in Section III.1, we can obtain all physical quantities of interest for this spacetime. Therefore, we have
@@
Figure 3: Scalar field Eq. (35) for some charge values and considering regions inside and outside any possible horizon is represented in a). Similarly, we have the potential Eq. (37) b) and the electromagnetic functions Eq. (36) c) and Eq. (38) d). All for $n=1/3$ setting, adopting $F_{0}=1$. The vertical lines represent the positions of the horizons. For this model, there are four horizons: two event horizons (the outer horizons) and two Cauchy horizons (the inner horizons). In the cases presented, horizons exist only for $q_{m}<4m/(3\sqrt{3})\approx 0.7698m$.
In Fig. 3, we have a graphical representation of the physical quantities obtained above for the configuration $n=1/3$ for Bardeen-type spacetime. In Fig. 3, we show the behavior of the scalar field, Eq. (35), in all regions of the spacetime, considering different charge values. For some values of charge, $q_{m}>q_{ext}=\frac{4m}{3\sqrt{3}}$, there is no event horizon. Note in the asymptotic form of the scalar field Eq. (41) that for a certain limiting value of charge within the horizon, around $q_{m}\approx{0.26m}$, the scalar field tends to reverse sign.
For the scalar potential, Fig. 3, we observe a behavior similar to the cases previously investigated. For regions outside the external horizon, the potential behaves like a barrier. For regions within the horizon, the potential starts to exhibit two symmetrical minima that tend to grow as the charge becomes more internal to extreme radius. The electromagnetic quantities Eq. (36) and Eq. (38) exhibit behavior similar to that presented in Section III.1 and can be visualized in Figs. 3 and 3.
@@
The asymptotic form for the scalar field Eq. (42) is written as:
$\phi\left(x\to\infty\right)=-\phi\left(x\to{-\infty}\right)=\frac{5\sqrt{5/2}\left(31185\pi m^{2}-524288mq_{m}+485100\pi q_{m}^{2}\right)}{88704q_{m}^{6}},$ (47)
where $F_{0}=1$.
Figure 4: Scalar field, Eq. (42), in panel a) and potential, Eq. (43), in panel b) with different charge values and considering regions inside and outside any possible horizon, for the configuration $n=1/5$ Here, we define $F_{0}=1$. The vertical lines represent the positions of the horizons. For this model, there are four horizons: two event horizons (the outer horizons) and two Cauchy horizons (the inner horizons). In the cases presented, horizons exist only for $q_{m}<4m/(3\sqrt{3})\approx 0.7698m$.
In Fig. 4, we have the graphical representation for the scalar field Eq. (42) and potential Eq. (43) with a configuration $n=1/5$. Qualitatively, the scalar field Fig. 4 and the potential Fig. 4 exhibit similar behavior to the previously investigated cases, with only minor numerical shifts. From the asymptotic form for the scalar field in Eq. (47), we find that there are no real values for the charge that allow a sign change in this expression, identical to what was observed in the Simpson-Visser case with $n=1/5$.
@@
The same analysis can be extended to the second model (Bardeen-type) when comparing the functions Eqs. (39) and (46) as well as the expression Eq. (36) which are the same found for the canonical scalar field [52].
@@ V General Energy Conditions
We know that black-bounce solutions violate the null energy condition, which must violate all other energy conditions [14, 53]. In this section, we will make the generic construction for the energy conditions, that is, just considering the static and spherically symmetric metric Eq. (14), as well as the stress-energy tensor defined in Eqs. (5-6). The energy conditions are defined by the combination of energy density and radial and tangential pressures as [^1] In the notation we follow, NEC is used for the null energy condition, WEC for the weak energy condition, SEC for the strong energy condition, and DEC for the dominant energy condition. The subscript $1$ is used when we have the energy density with radial pressure, $2$ when we have the energy density with tangential pressure and $3$ when we have both pressures or just the energy density.
@@
where the energy density is denoted by $\rho^{\phi,EM}$, while $p^{\phi,EM}_{1}$ and $p^{\phi,EM}_{2}$ represent the radial and tangential pressures, respectively. This expression for the stress-energy tensor above, Eq. (52), refers to the part outside the event horizon, $A>0$. Within the event horizon, $A<0$, we have: